OpenStudy (anonymous):

Consider a balloon that has a volume V. It contains n moles of gas, it has an internal pressure of P, and its temperature is T. If the balloon is heated to a temperature of 15.5T while it is placed under a high pressure of 15.5P, how does the volume of the balloon change? ~ It doubles. ~ It stays the same. ~ It increases greatly. ~ It decreases slightly.

OpenStudy (anonymous):

@DanJS You wouldn't happen to know how to solve this would you? (sorry for the constant nagging. It's my last day of chemistry and I'm just trying to get it all done.)

OpenStudy (anonymous):

@jim_thompson5910 Are you good with chemistry?

OpenStudy (anonymous):

@ganeshie8 Could you help me by chance?

OpenStudy (anonymous):

@TheSmartOne @iGreen Help please?

OpenStudy (unklerhaukus):

The ideal gas equation \[PV=nRT\] The number of moles of gas in the ballon a constant \[\frac{PV}{T}=nR\] So for any different pressures, volumes, temperatures \[\frac{P_1V_1}{T_1}=\frac{P_2V_2}{T_2}\]

OpenStudy (unklerhaukus):

\[T_2=15.5\,T_1\] \[P_2=15.5\,P_1\] solve for \(V_2\), in terms of \(V_1\)